OpenStudy (anonymous):

Given the linear equation y = -4x - 2, find the y-coordinates of the points (1,),(-1,), and (-3,).

OpenStudy (anonymous):

y=-2(2x+1)

OpenStudy (anonymous):

\[y=-2(2x+1)\]

OpenStudy (anonymous):

y =-4x-2 so it can be written as y= y=−2(2x+1)